OpenStudy (anonymous):

URGENT HELP. The height of a triangle is 5 m less than its base. The area of the triangle is 42 m2. Find the length of the base.

OpenStudy (anonymous):

The altitude of a triangle is 5 m less than its base. The area of the triangle is 42 m2. Find the length of the base. . Let b = base then b-5 = altitude . since we know that the area for any triangle is (1/2)bh where b is base h is height or altitude . (1/2)b(b-5) = 42 b(b-5) = 84 b^2-5b = 84 b^2-5b-84 = 0 (b-12)(b+7) = 0 b = {12, -7} toss out the negative solution leaving: base = 12 meters